Two, long, straight wires separated by 0.40 cm car
Answer #1

1) The attraction force is given by F = i (l x B) where length vector is taken in direction of current and B is magnetic field. Magnetic field is given by

B= u*i/2piR with direction given by l x r where l is length of Magnetic field causing wire and r is distance vector where field has to be found out. For wire 2 the magnetic field experienced by it due to wire 1 is in direction given by l x r which is going inside the paper direction. Hence force direction is cross product of upward direction and inside the paper direction which comes out to be right sided attraction force. Hence wires are attracted.

2) At P due to wire 1 = u*i/2pi*7 + u*i/2pi*4 = 11u*i/ 28*2pi = 11u*i / (56pi) where u = 4*pi*10^7

Direction is inwards as discussed above.